OpenStudy (anonymous):

Find values of a and b such that a is a whole number, b is a rational number but not an integer, and a divided by b equals -8

OpenStudy (anonymous):

It doesnt have to be ONE number it can be whatever fits the equation.

OpenStudy (anonymous):

\[\frac{a}{b}=-8\]But b is rational so we can write it as a ratio of two intergers c and d:\[\frac{a}{\frac{c}{d}}=-8\]Thus\[ad=-8c\]You're just asked to find values which satisfy this. Picking a=4, c=2 then d=-4. So b=-2/4 of -1/2. The solution is not unique.

OpenStudy (anonymous):

typo...should read b=-2/4 OR -1/2

OpenStudy (anonymous):

Haha thanks so much!! Your a life saver!

OpenStudy (anonymous):

no prob :)

OpenStudy (anonymous):

but a has to be a whole number, that doesnt work

OpenStudy (anonymous):

Wait nevermind!